OpenStudy (anonymous):

A ball is hurled in the air such that its horizontal (x) and vertical (y) shifts are given by the functions x(t) = 20t and y(t) = 40t − 5t2, where x and y are in meters and the time t is in seconds. What is the ratio of y to x at time t = 5 seconds?

OpenStudy (anonymous):

0.25 0.50 0.75 0.85

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

\[\frac{y}{x}\] evaluate this at t=5

OpenStudy (anonymous):

40(5)-5(5)^2/20(5)

OpenStudy (freckles):

\[\frac{40(5)-5(5)^2}{20 \cdot 5}\] yes evaluate that :)

OpenStudy (anonymous):

0.75?

OpenStudy (anonymous):

you know what math im doing?

OpenStudy (anonymous):

@freckles